# Quantum in Biology, Quantum for Biology, and Biology for Quantum: Mapping the Evidence and the Road Ahead

### 30 Apr 2026 v1

@@ Appendix A Summary tables across the map
Topic Central claim Strongest evidence base Main confounds Decisive tests DNA replication Tautomerization and mispairing may be influenced by proton tunneling or proton delocalization [138, 183] Quantum-chemistry plausibility and mechanistic modeling [183, 206] Polymerase conformational selection, proofreading, and classical fluctuations Isotope and temperature scaling in polymerase-resolved assays with tautomer-lifetime and misincorporation constraints Enzyme catalysis Hydrogen transfer includes tunneling contributions [116] Kinetic isotope effects, temperature trends, pressure or viscosity perturbations, and constrained theory Coupled catalytic steps, gating motions, and shifting rate limitation Targeted substitution or directed-evolution series with constrained modeling and orthogonal structural constraints [119] PCET motifs Coupled electron–proton transfer governs key redox and proton-transfer steps [149, 86] Physical-chemistry theory and targeted model systems Attribution within full catalytic cycles and competition from semiclassical rate descriptions Isotope, pH, and electrostatic scaling in isolated biological steps while holding the redox landscape as constant as possible Vision Near-quantum-limited photon detection and ultrafast photochemical dynamics in rhodopsin [91, 185] Single-photon physiology, ultrafast photochemistry, and quantum-resource-theoretic analyses Coherence signatures can be mimicked by vibrational wavepackets, ensemble effects, and excitation-condition dependence Controlled dephasing or vibronic perturbations linked to quantum yield, isomerization timing, or downstream amplification statistics Olfaction Vibrationally assisted electron transfer may contribute to receptor activation [28, 207, 220] Theoretical and open-system plausibility under receptor-like dissipative conditions [220, 47] Binding and dynamical explanations; limited receptor-level evidence [29, 94] Receptor-resolved isotopologue tests controlling affinity and binding pose, together with detection of charge-transfer signatures Magnetoreception Radical-pair spin dynamics can yield magnetic sensitivity [186, 97, 41] Open-system spin models, chemically constrained estimates, and proxy spin-chemistry experiments [5, 169] Species dependence, competing mechanisms, radical identity, and transduction uncertainty [236] Genetics, spectroscopy, identified radicals, and controlled electromagnetic perturbations tied to predicted behavioral or physiological effects Photosynthesis Coherence-like or vibronically mixed excitonic dynamics may influence energy transport [66, 166, 54] Reproducible ultrafast signatures and open-system theory [100, 49] Vibrational mimicry, ensemble effects, static disorder, and difficulty proving in vivo functional advantage [189] Selective detuning or dephasing perturbations linked to functional yield under near-native conditions DNA charge transport DNA-mediated redox transport may contribute to lesion sensing or redox signaling [150, 208] Physicochemical evidence and repair-protein redox studies [150, 208] Competing protein recruitment, diffusive search, chromatin context, and lesion-recognition steps may dominate repair kinetics Nucleosome- or chromatin-like substrate tests and cell-based assays with redox mutants plus quantitative model comparison DNA repair photochemistry Photolyase uses excited-state photochemistry and ultrafast electron transfer to repair lesions [191, 217] Mechanistic photochemistry, identified cofactors, and repair kinetics Stronger coherence claims remain difficult to isolate in fluctuating protein environments and depend on excitation conditions Perturbations that selectively alter dephasing or related excited-state dynamics and test corresponding changes in repair yield or pathway branching Cellular respiration Tunneling corrections may influence electron or proton transfer in respiratory complexes [156, 89] Mechanistic modeling, transfer-pathway calculations, and biochemical constraints Classical transfer with gating, reorganization, membrane-potential effects, and redox-landscape changes Perturbations isolating tunneling pathways with orthogonal biochemical validation and controlled redox conditions CISS Chiral biomolecular scaffolds could influence spin selectivity in electron transfer and downstream spin-sensitive chemistry [30, 8] Strong interface and junction evidence for CISS itself [30, 8] Interfacial artifacts, contact geometry, and classical parameter changes Solution-phase or near-physiological tests with chirality-matched controls and independent spin-sensitive readouts Ion channels Coherent-transport hypotheses in selectivity filters could contribute to selectivity or conduction under some conditions [224, 75] Theoretical proposals; established quantum chemistry of ion energetics and proton pathways [188, 60] Classical stochastic models reproduce many conductance and selectivity observables Perturbations that isolate coherence or resonance predictions without broadly shifting classical energetics Microtubules Excited-state transport, collective optical effects, or spin-sensitive signatures may occur in tubulin assemblies [108, 56] In vitro signatures, magnetic-isotope perturbation reports, and multiple photophysical models [107, 168] Preparation artifacts, sample dependence, competing quenching pathways, and structured classical hopping Independent replication with standardized samples and discriminating observables under physiologically relevant or explicitly controlled in vitro conditions Genetic code and quantum formalisms Quantum formalisms may describe codon structure or coding patterns, but do not by themselves establish physical quantum processing in living systems [70, 24, 167, 18] Descriptive mathematical analyses and speculative origin-of-coding frameworks [70, 24, 18] Flexible formalisms; evolutionary, stereochemical, and error-tolerance explanations already account for many code features [81, 233, 232] Either a measurable quantum signature affecting coding or regulation, or a predictive advantage under strict model comparison and out-of-sample validation Posner clusters Protected nuclear spins in Ca9(PO4)6 may influence biochemistry [71] Structural plausibility, computational studies, and lifetime constraints [212, 176] Formation, symmetry, spin relaxation, entanglement generation, and biochemical transduction under physiological conditions In situ detection plus spin-relaxation, coherence, and nonclassical-correlation measurements under physiologically relevant conditions
@@
Topic Central idea Strongest evidence base Main uncertainty Benchmarks DNA origami scaffolds Programmable self-assembly positions emitters and nanostructures with nanometer-scale precision [121, 128, 96, 241] Robust nanoscale assembly demonstrations Stability, yield, and integration in device settings Reproducible coupling distributions and integration with photonics Virus-templated materials Biological templates guide synthesis and assembly of nanowires and related nanostructures [141] Controlled synthesis and assembly in vitro Whether loss and disorder can meet quantum-device requirements Performance distributions benchmarked against top-down fabrication DNA-templated wiring Biological scaffolds define conductive nanoscale pathways [35] Electrically connected nanoscale structures Interface quality and purity under quantum-device conditions Loss, noise, and reproducibility under device conditions Biohybrid polaritons Biological chromophore networks can strongly couple to optical cavities [53] Demonstrated strong coupling in biohybrid systems Stability, controllability, and scalability Reproducible coupling, tunability, and circuit integration Genetically encoded qubits Biomolecules can host controllable quantum states [68] Demonstrated qubit platform in a protein Fidelity and coherence in complex environments Standard qubit metrics and reproducibility across expression contexts Bioinspired transport Noise-assisted transport and interference-inspired design principles motivated by photosynthesis [177, 57] Theory and early prototypes [187] Robustness to disorder and scalability Device gains that persist across disorder ensembles Brain-inspired quantum and quantum-inspired neural computation Brain-inspired architectures motivate both quantum neuromorphic designs and quantum-like models for complex data, cognition, and neuroimaging [126, 144, 34, 3] Classical neuromorphic hardware, proof-of-principle quantum neuromorphic proposals, and quantum-like modeling frameworks [126, 34, 198, 111] Whether such approaches outperform strong classical neuromorphic, deep-learning, or standard quantum baselines on well-defined tasks Matched benchmarks for scalability, robustness, resource demands, and generalization against strong classical and standard quantum baselines

### 27 May 2026 v2

@@ Appendix A Summary tables across the map
Topic Central claim Strongest evidence base Main confounds Decisive tests DNA replication Tautomerization and mispairing may be influenced by proton tunneling or proton delocalization [147, 197] Quantum-chemistry plausibility and mechanistic modeling [197, 222] Polymerase conformational selection, proofreading, and classical fluctuations Isotope and temperature scaling in polymerase-resolved assays with tautomer-lifetime and misincorporation constraints Enzyme catalysis Hydrogen transfer includes tunneling contributions [121] Kinetic isotope effects, temperature trends, pressure or viscosity perturbations, and constrained theory Coupled catalytic steps, gating motions, and shifting rate limitation Targeted substitution or directed-evolution series with constrained modeling and orthogonal structural constraints [125] PCET motifs Coupled electron–proton transfer governs key redox and proton-transfer steps [158, 91] Physical-chemistry theory and targeted model systems Attribution within full catalytic cycles and competition from semiclassical rate descriptions Isotope, pH, and electrostatic scaling in isolated biological steps while holding the redox landscape as constant as possible Vision Near-quantum-limited photon detection and ultrafast photochemical dynamics in rhodopsin [96, 199] Single-photon physiology, ultrafast photochemistry, and quantum-resource-theoretic analyses Coherence signatures can be mimicked by vibrational wavepackets, ensemble effects, and excitation-condition dependence Controlled dephasing or vibronic perturbations linked to quantum yield, isomerization timing, or downstream amplification statistics Olfaction Vibrationally assisted electron transfer may contribute to receptor activation [33, 224, 237] Theoretical and open-system plausibility under receptor-like dissipative conditions [237, 52] Binding and dynamical explanations; limited receptor-level evidence [34, 99] Receptor-resolved isotopologue tests controlling affinity and binding pose, together with detection of charge-transfer signatures Magnetoreception Radical-pair spin dynamics can yield magnetic sensitivity [200, 102, 46] Open-system spin models, chemically constrained estimates, and proxy spin-chemistry experiments [5, 181] Species dependence, competing mechanisms, radical identity, and transduction uncertainty [257] Genetics, spectroscopy, identified radicals, and controlled electromagnetic perturbations tied to predicted behavioral or physiological effects Photosynthesis Coherence-like or vibronically mixed excitonic dynamics may influence energy transport [72, 178, 60] Reproducible ultrafast signatures and open-system theory [105, 54] Vibrational mimicry, ensemble effects, static disorder, and difficulty proving in vivo functional advantage [205] Selective detuning or dephasing perturbations linked to functional yield under near-native conditions DNA charge transport DNA-mediated redox transport may contribute to lesion sensing or redox signaling [159, 225] Physicochemical evidence and repair-protein redox studies [159, 225] Competing protein recruitment, diffusive search, chromatin context, and lesion-recognition steps may dominate repair kinetics Nucleosome- or chromatin-like substrate tests and cell-based assays with redox mutants plus quantitative model comparison DNA repair photochemistry Photolyase uses excited-state photochemistry and ultrafast electron transfer to repair lesions [207, 234] Mechanistic photochemistry, identified cofactors, and repair kinetics Stronger coherence claims remain difficult to isolate in fluctuating protein environments and depend on excitation conditions Perturbations that selectively alter dephasing or related excited-state dynamics and test corresponding changes in repair yield or pathway branching Cellular respiration Tunneling corrections may influence electron or proton transfer in respiratory complexes [165, 94] Mechanistic modeling, transfer-pathway calculations, and biochemical constraints Classical transfer with gating, reorganization, membrane-potential effects, and redox-landscape changes Perturbations isolating tunneling pathways with orthogonal biochemical validation and controlled redox conditions Ultra-weak photon emission Biological systems emit ultra-weak photons linked to oxidative metabolism, ROS chemistry, electronically excited species, and mitochondrial redox activity [192, 167, 243] Recent studies support UPE as a possible marker of biological state and selected radiation-bystander or mitochondrial responses [168, 29, 138, 240, 166] Photon emission alone does not demonstrate coherence, nonclassical light, functional optical communication, or source-specific metabolic inference [57] Calibrated spectral, temporal, and photon-statistical measurements with independent ROS, metabolic, electrophysiological, optical-transport, and tissue-calibration controls CISS Chiral biomolecular scaffolds could influence spin selectivity in electron transfer and downstream spin-sensitive chemistry [35, 8] Strong interface and junction evidence for CISS itself [35, 8] Interfacial artifacts, contact geometry, and classical parameter changes Solution-phase or near-physiological tests with chirality-matched controls and independent spin-sensitive readouts Ferritin protein nanoparticles Ferritin cores and protein shells may support electron tunneling, Coulomb-blockade-like behavior, and magnetic effects relevant to bioelectric or biomagnetic hypotheses [67, 203] Ferritin junction and multilayer transport experiments, together with review-level synthesis of ferritin electrical and magnetic properties [134, 32, 202] Device geometry, iron loading, hydration, aggregation, redox state, cellular localization, and uncertain linkage to biological function In situ or near-native perturbation tests controlling iron loading, redox state, magnetic field, chirality or spin-sensitive readouts, and defined biological endpoints Ion channels Coherent-transport hypotheses in selectivity filters could contribute to selectivity or conduction under some conditions [244, 81] Theoretical proposals; established quantum chemistry of ion energetics and proton pathways [204, 65] Classical stochastic models reproduce many conductance and selectivity observables Perturbations that isolate coherence or resonance predictions without broadly shifting classical energetics Microtubules Excited-state transport, collective optical effects, or spin-sensitive signatures may occur in tubulin assemblies [113, 62] In vitro signatures, magnetic-isotope perturbation reports, and multiple photophysical models [112, 180] Preparation artifacts, sample dependence, competing quenching pathways, and structured classical hopping Independent replication with standardized samples and discriminating observables under physiologically relevant or explicitly controlled in vitro conditions Genetic code and quantum formalisms Quantum formalisms may describe codon structure or coding patterns, but do not by themselves establish physical quantum processing in living systems [76, 27, 179, 21] Descriptive mathematical analyses and speculative origin-of-coding frameworks [76, 27, 21] Flexible formalisms; evolutionary, stereochemical, and error-tolerance explanations already account for many code features [87, 254, 253] Either a measurable quantum signature affecting coding or regulation, or a predictive advantage under strict model comparison and out-of-sample validation Posner clusters Protected nuclear spins in Ca9(PO4)6 may influence biochemistry [77] Structural plausibility, computational studies, and lifetime constraints [229, 189] Formation, symmetry, spin relaxation, entanglement generation, and biochemical transduction under physiological conditions In situ detection plus spin-relaxation, coherence, and nonclassical-correlation measurements under physiologically relevant conditions
@@
Topic Central idea Strongest evidence base Main uncertainty Benchmarks DNA origami scaffolds Programmable self-assembly positions emitters and nanostructures with nanometer-scale precision [127, 135, 101, 262] Robust nanoscale assembly demonstrations Stability, yield, and integration in device settings Reproducible coupling distributions and integration with photonics Virus-templated materials Biological templates guide synthesis and assembly of nanowires and related nanostructures [150] Controlled synthesis and assembly in vitro Whether loss and disorder can meet quantum-device requirements Performance distributions benchmarked against top-down fabrication DNA-templated wiring Biological scaffolds define conductive nanoscale pathways [40] Electrically connected nanoscale structures Interface quality and purity under quantum-device conditions Loss, noise, and reproducibility under device conditions Biohybrid polaritons Biological chromophore networks can strongly couple to optical cavities [59] Demonstrated strong coupling in biohybrid systems Stability, controllability, and scalability Reproducible coupling, tunability, and circuit integration Genetically encoded qubits Biomolecules can host controllable quantum states [74] Demonstrated qubit platform in a protein Fidelity and coherence in complex environments Standard qubit metrics and reproducibility across expression contexts Bioinspired transport Noise-assisted transport and interference-inspired design principles motivated by photosynthesis [190, 63] Theory and early prototypes [201] Robustness to disorder and scalability Device gains that persist across disorder ensembles Quantum biomimetics and artificial-life-inspired algorithms Life-like behaviors such as self-replication, mutation, interaction, inheritance, adaptation, and death can be used as design motifs for quantum-information protocols and controllable artificial-life models [12, 13, 242] Quantum artificial-life models, biomimetic cloning protocols, and proof-of-principle implementation on an IBM quantum computer [11, 13, 136] Currently exploratory; unclear advantage over classical artificial-life models, evolutionary algorithms, or standard quantum algorithms Matched benchmarks testing scalability, resource use, robustness, and task-level advantage against classical artificial-life and standard quantum baselines Brain-inspired quantum and quantum-inspired neural computation Brain-inspired architectures motivate both quantum neuromorphic designs and quantum-like models for complex data, cognition, and neuroimaging [132, 153, 39, 3] Classical neuromorphic hardware, proof-of-principle quantum neuromorphic proposals, and quantum-like modeling frameworks [132, 39, 214, 116] Whether such approaches outperform strong classical neuromorphic, deep-learning, or standard quantum baselines on well-defined tasks Matched benchmarks for scalability, robustness, resource demands, and generalization against strong classical and standard quantum baselines